# A Hierarchy of Topological and Superconducting States in Rhombohedral Hexalayer Graphene

### v1

@@ A Hierarchy of Topological and Superconducting States in Rhombohedral Hexalayer Graphene
Having considered two possible scenarios to account for our observations, we note that alternative explanations may still be viable. Nonetheless, the discovery of a superconducting phase that is stabilized—rather than suppressed—by an out-of-plane magnetic field, and its striking connection to an exotic sequence of quantum Hall states, represents a fundamentally new phenomenon. The intriguing interplay between superconductivity and quantum Hall effect presents a compelling challenge to existing theoretical paradigms, which is expected to stimulate broad future interest into emergent phenomena of strongly interacting electrons in 2D material heterostructures.

### v2

@@ A Hierarchy of Topological and Superconducting States in Rhombohedral Hexalayer Graphene
Besides the CDW scenario, we note that alternative explanations may still be viable. Nonetheless, the discovery of a superconducting phase that is stabilized—rather than suppressed—by an out-of-plane magnetic field, and its striking connection to an exotic sequence of quantum Hall states, represents a fundamentally new phenomenon. The intriguing interplay between superconductivity and quantum Hall effect presents a compelling challenge to existing theoretical paradigms, which is expected to stimulate broad future interest into emergent phenomena of strongly interacting electrons in 2D material heterostructures.